OpenStudy (anonymous):

A university is conducting a survey to study the number of hours students sleep at night during exam week. A sample of 20 students was drawn, and these responses were obtained: 6.5, 5, 8, 5.5, 7, 4, 4.5, 5, 6, 5.5, 4, 7.5, 6, 7, 8, 5, 6, 6.5, 6.5, 5. The standard deviation of the population mean is 2 hours. Arrange the steps in the correct sequence to find the range in which 95% of the sample mean occurs.

OpenStudy (anonymous):

occurs. STEPS: The range in which 95% of the sample mean occurs = 5.93 + 2 x 0.45 and 5.93 - 2 x 0.45 = 6.83 hours and 5.03 hours. The sample mean is u = 118.5/20 = 5.93. Calculate the standard deviation. The 20 students' total number of study hours is 118.5 hours. The standard error of the mean of the sample is o/sqrtN = 2/sqrt20 = 0.45 The range in which 95% of the sample mean occurs is u +- 2 x o/sqrtN

OpenStudy (odotellessa):

wow they sleep little >_> there's only one two 8s... seems like they haven't been educated enough about the benefits of sleep enough apparently.